OpenStudy (anonymous):

given that 5 exponent -m is equal to x..find an expression in x, for 5 exponent 1+m..

OpenStudy (anonymous):

\[5^{-m}\]

OpenStudy (anonymous):

= x right

OpenStudy (anonymous):

we have to find \[5^{m+1}\]

OpenStudy (anonymous):

ok go on

OpenStudy (anonymous):

ok see , 5^(-m) = x we have to find 5^(m+1) according to laws of exponents a^(b+c) = a^b * a^c hence 5^(m+1) = 5^m * 5^1 = 5^m * 5 1/(5^m) = x 1/x = 5^m 5^m * 5 = 5^(m+1) 1/x * 5 = 5^(m+1) 5/x = 5^(m+1)

OpenStudy (anonymous):

got it now agradecompetent

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

any other question ?

OpenStudy (anonymous):

no i solved all or u wanna try them?

OpenStudy (anonymous):

\[\color{indigo}{\text{yes of course i would try them :) may be i would be successful in them or not. let us see }}\]

OpenStudy (anonymous):

given that {log base 8 (p+2)+log base 8q}=r and {log base 2 (p-2)-log base 2q=2r+1, show that p^2=4+2*32r

OpenStudy (anonymous):

it is rather show that p^2=4+2*32^r